integrate sqrt(cosx)/sinx

Grade:11

3 Answers

Parvez ali
10 years ago
integration(cosx/sinx.sqrt(1-sin2x))dx
let sinx=t, cosx dx =dt
=integration(1/t.sqrt(1-t2))dt
let t=1/u, dt=-1/u2 du
=integration(-1/sqrt(u2-1))du
=-log(u+sqrt(u2-1))+c
=-log(cosecx+sqrt(cot2x))+c
=-log(cosecx+modulus(cotx )+c
Thanks & Regards
Parvez Ali
askIITians Faculty
B-tech
ISM Dhanbad
Priya Pankaj
8 years ago
how come sqrt(cosx)=sqrt(1-sin^2x).I think the que is integrate sqrt(cos2x)/sinx.then it would be cos2x/{sinx.(cos2x)}
Priya Pankaj
8 years ago
how come sqrt(cosx)=sqrt(1-sin^2x).I think the que is integrate sqrt(cos2x)/sinx.then it would be cos2x/{sinx.(cos2x)^1/2}